Test the uniform convergence of the series

Σ∞

𝑛=1 x[𝑛/(1+𝑛^2π‘₯^2)βˆ’(𝑛+1)/(1+(𝑛+1)^2π‘₯^2)]


1
Expert's answer
2021-07-16T02:41:38-0400

"Σ^∞_{𝑛=1} x[\frac{𝑛}{(1+𝑛^2𝑥^2)−\frac{(𝑛+1)}{(1+(𝑛+1)^2𝑥^2)}}]"

We can reduce the series to a simpler form to work with as shown below

"Σ^∞_{𝑛=1} [\frac{x}{(1+𝑛𝑥)((n-1)x+1)}]"

The above series is uniformly convergent on any finite interval [a; b], when "0 < a < b" . But the series is only point wise convergent but not uniformly convergent on [a, b]

Here "U_n(x)= [\frac{x}{(1+𝑛𝑥)((n-1)x+1)}] = \frac{1}{(n-1)x+1}-\frac{1}{nx+1}"

Therefore "s_n(x) = u_1(x) + u_2(x) + u_3(x) + \u00b7 \u00b7 \u00b7 + u_n(x)"

"s_n(x)= (1-\frac{1}{x+1})+(\frac{1}{x+1}+\frac{1}{2x+1})+(\frac{1}{2x+1}+\frac{1}{3x+1})+...+[\frac{1}{(n-1)x+1}-\frac{1}{nx+1}]\\
s_n(x)= 1-\frac{1}{nx+1}\\
\therefore s(x)=Lim_{n \to \infin} s_n(x)= x = \begin{cases}
 1 &\text{if } x\not=0 \\
 0 &\text{if } x=0
\end{cases}"

Therefore the sum function s(x) is discontinuous on [a, b], and therefore the convergence is not uniform

on [a, b], it is only point wise. When "x \not= 0" , let "\epsilon > 0" be given

"|s_n(x)-s(x)|=|1-\frac{1}{nx+1}-1|=|\frac{1}{nx+1}| < \epsilon" ; whenever "n > \frac{1}{x}(\frac{1}{\epsilon}-1)"

But "\frac{1}{x}(\frac{1}{\epsilon}-1)" decreases with x. Hence if we takeΒ  "N= [\max\limits_{x \epsilon[a,b] } \begin{cases}
 
\end{cases} \frac{1}{x}(\frac{1}{\epsilon}-1)] +1\epsilon N"

which is independent of x, then we obtain "|s_n(x)-s(x)|< \epsilon" whenever n> N whenever n > N for all "x \epsilon [a, b]" i.e. the series converges uniformly to s(x) = 1 on [a,b].